OpenStudy (anonymous):

HELP!!!!!!!!!!!!!!!!!!! A system of equations is shown below: 8x+5y=9 3x+y=4 Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. show the steps to do this.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

ok... so multiply the 2nd equation by - 5 and add it to the 1st equation... 8x + 5y = 9 -15x - 5y = -20 ---------------- -7x = -11

OpenStudy (anonymous):

im sorry what did you do?

OpenStudy (campbell_st):

the 2nd equation is 3x + y = 4 multiply every term by -5 -15x -5y = -20 then add this new equation to the 1st equation -15x - 5y + 8x + 5y = -20 + 9 -8x = -11

OpenStudy (anonymous):

but don't I have to create a system of equations ?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

-8x=11 will be my answer?

OpenStudy (anonymous):

part b: show that the equivalent system has the same solution as the original system of equations

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

part b: show that the equivalent system has the same solution as the original system of equations

OpenStudy (anonymous):

part b: show that the equivalent system has the same solution as the original system of equations

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

PLEASE HELP IT IS MY LAST QUESTION AND IM REALLY EXHAUSTED

hero (hero):

8x+5y=9 3x+y=4 Multiply the second equation by -4, then add the first equation to it: -4(3x + y) = -4(4) -12x - 4y = -16 -12x + 8x + 5y - 4y = 9 - 16 -4x + y = -7 So the equivalent system is 8x + 5y = 9 -4x + y = -7

hero (hero):

Solve one of the systems to get (x, y) value, then test the value in the other system.

OpenStudy (anonymous):

THANK YOU!!

OpenStudy (anonymous):

im stuck on that part finding the x and y values

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

can you please help?

OpenStudy (anonymous):

8x+5y=9 3x+y=4 Multiply the second equation by -4, then add the first equation to it: -4(3x + y) = -4(4) -12x - 4y = -16 -12x + 8x + 5y - 4y = 9 - 16 -4x + y = -7 So the equivalent system is 8x + 5y = 9 -4x + y = -7

OpenStudy (anonymous):

then I have to solve show that the equivalent system has the same solution as the original system of equations

OpenStudy (nikato):

lets start with your original 8x+5y=9 3x+y=4 1/multiply the second by -5 2. add the first equation and the new second equation up. *the y's should be able to cancel each other 3. you should be able to solve for x now

OpenStudy (anonymous):

x=11/7

OpenStudy (anonymous):

y=-5/7

OpenStudy (nikato):

yes so (11/7 , -5/7) is the solution for this

OpenStudy (nikato):

now do the same with the equivalent equation 8x+5y=9 -4x+y=-7 1. multiply second equation by 2 2. add the two equation up. 3. you should be able to solve for y

OpenStudy (anonymous):

thnk you !!!